Someone please help me with this question point b)

a) Will this change in capital structure affect the market value of the firm? Discuss

b) How will the return on equity be affected by this change? What is the new return on equity of the company?

Fan Plc is a publicly traded firm. The market value of its equity is £70,000,000 and its debt £30,000,000. The yield to maturity of the debt is 5%, the shareholders require a 20% return, and the company pays 30% corporate tax. They have recently decided to repurchase £10,000,000 worth of equity, and finance the repurchase through the issuance of new debt.
a) Will this change in capital structure affect the market value of the firm? Discuss
b) How will the return on equity be affected by this change? What is the new return on equity of the company?
Answer #1

Yes, the market value of firm changes. the market value of firm is generally calculated as market value of equity which is 70 - 10 in the post repurchase scenario. the increase debt has changed the interest expenses which inturn increased the tax savings, hence the required return on equity increases.

the effect of market value change in equity will change the rate of required return for equity from 20% to 22.5%.

D/E before repurchase = 30/70 (in Millions)

D/V before repurchase = 30%

E/V before repurchase = 70%

Weight average cost of capital (WACC) = Re * We + Wd*Rd*(1-t)

Whereas re = Cost of equity

We = E/V

Rd = cost of debt

Wd = D/V

t = tax rate

given, Re = 20%

                Rd = 5%

                T = 30%

WACC = 0.7*0.2 + 0.3* 0.05 *(1-0.3)

                =0.1505 = 15.05%

Post Repurchase Scenario:

Equity value = 70-10 = 60

Debt value = 30 + 10 = 40

New D/V = 40/100 = 40%

New E/V = 60/100 = 60%

New Re = WACC – (Rd* Wd*(1-t))/(We)

                = (0.1505-(0.4*0.05*(1-0.3))/(0.6)

                =0.2275 = 22.75%